OpenStudy (matlee):

A number N is multiplied by 3. The result is the same as when N is divided by 3. What is the value of N?

ILovePuppiesLol (ilovepuppieslol):

well i mean

OpenStudy (matlee):

Help

OpenStudy (matlee):

i know the naswer is 0 but i dont know the work

ILovePuppiesLol (ilovepuppieslol):

0 times 3 is 0 and 0 divided by 3 is 0

OpenStudy (matlee):

With work i got 9

ILovePuppiesLol (ilovepuppieslol):

n(3)= n/3 ? or something like that

OpenStudy (matlee):

i got 9 for N

ILovePuppiesLol (ilovepuppieslol):

9 times 3 is 27 and 9 divided by 3 is 3

ILovePuppiesLol (ilovepuppieslol):

those are not equal

ILovePuppiesLol (ilovepuppieslol):

show me ur work

OpenStudy (matlee):

n(3)=n/3 so multiply by 3 to take out the division. (3)n(3)=n/(3*3) then 9N=N so divide N and get 9=1?

ILovePuppiesLol (ilovepuppieslol):

u don't get that, u just figure it out by substituting 0 for n, that is the only number that makes sense

ILovePuppiesLol (ilovepuppieslol):

does it ask you to show work?

OpenStudy (matlee):

Yes

ILovePuppiesLol (ilovepuppieslol):

well I'm pretty sure u are overthinking this

ILovePuppiesLol (ilovepuppieslol):

just do it the way i showed u, it should work out that way

zepdrix (zepdrix):

Hey Mat, if you would like some algebraic steps as to how to solve for N, then here you go,\[\large\rm 3N=\frac{N}{3}\]Multiply through by 3,\[\large\rm 9N=N\]Subtract N to the other side,\[\large\rm 9N-N=0\]Combine like-terms,\[\large\rm 8N=0\]Divide by 8,\[\large\rm N=0\] With equations, dividing out variables can sometimes lead to big problems. So it's better to take a more length route like this. You'll notice that if you chose to divide both sides by N, there is no longer an N in your equation to solve for.